OpenStudy (thecatman):

A waiter earned $16 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 12 bills. 1 2 3 4

OpenStudy (anonymous):

Best AnswerVoter's Choice D.W. answered 3 years ago "$16 in $1 bills and $5 bills." x = number of one-dollar bills y = number of five-dollar bills 1x + 5y = 16 "he received a total of 12 bills" x + y = 12 x = 12 - y 1(12-y) + 5y = 16 12 - y + 5y = 16 4y = 4 y = 1 x = 12 - y = 11 He received 1 five-dollar bill and 11 one-dollar bills Got this off Yahoo, might help :)

OpenStudy (thecatman):

thanks for helping

OpenStudy (anonymous):

You're welcome :)